OpenStudy (anonymous):

2 square root 27-3 square root 48+square root 147

OpenStudy (anonymous):

\[2\sqrt{27}-3\sqrt{48}+\sqrt{147}\]

OpenStudy (anonymous):

see 27=3^3 so you can write sqrt(27)=3*sqrt(3), if you write all the terms in this manner you can find the answer

OpenStudy (anonymous):

?

OpenStudy (anonymous):

step by step please

OpenStudy (anonymous):

27=3^3, so you can say that sqrt(27)=sqrt(3^3)=sqrt{(3^2)*(3)}=sqrt(3^2)*sqrt(3)=3*sqrt(3), you can change the other terms in this way and then you will have an equation like that a*sqrt(3)-b*sqrt(3)+c*sqrt(3)=(a-b+c)*sqrt(3)

OpenStudy (anonymous):

ok I am sorry

OpenStudy (anonymous):

lets start with 48, are you agree that 48=3*4*4

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so sqrt(48)=sqrt(3*4*4)=sqrt(3)*sqrt(4*4)

OpenStudy (anonymous):

so sqrt(48)=4*sqrt(3)

OpenStudy (anonymous):

allright

OpenStudy (anonymous):

confused about 147 how to i break it up

OpenStudy (anonymous):

if you look at the problem you can find that you need sqrt(3) in all the terms, so first divide 147 to 3 and it becomes 49

OpenStudy (anonymous):

you can write sqrt(147)=sqrt(3*49)=sqrt(3)*sqrt(49)=7sqrt(3)

OpenStudy (anonymous):

k so i get\[18\sqrt{3}-12\sqrt{3}+7\sqrt{3}\]

OpenStudy (anonymous):

no that is not true

OpenStudy (anonymous):

ya figured it out it \[6\sqrt{3}-12\sqrt{3}+7\sqrt{3}\]

OpenStudy (anonymous):

yes you are right, good luck.

OpenStudy (anonymous):

thank